a=2i+3j-k,
b=i-j+2k and
c=3i+4j+k.
find a+2b-c

Answer by greenestamps(13200)



Treat the vector components just like variables.

a=2i+3j-k,
b=i-j+2k,
c=3i+4j+k.

a+2b-c =
(2i+3j-k)+2(i-j+2k)-(3i+4j+k) =
(2+2-3)i + (3-2-4)j + (-1+4-1)k =
i-3j+2k